# What Language Models Know But Don’t Say: Non-Generative Prior Extraction for Generalization

### v1

@@ 4.2 LLM-Based Prior Extraction
We use the Gemma-2-27B, Gemma-3-27B, llama-3-8B, and GPT-2 large language models to extract logit-based priors as described in Section 3.2. We use these models in zero-shot mode without any fine-tuning or in-context examples. In our experiments, the queried lexical items are represented as single tokens across models, enabling direct logit-based comparisons.
Each dataset is processed independently. For each feature, we issue 10 prompts to extract token-level log-probabilities for the positive and negative completions. The number of prompts was selected via hyperparameter tuning (Section A.2.4). The resulting feature-wise prior means and variances are then used to define informative Gaussian priors in a Bayesian logistic regression model. Posterior inference is performed using PyMC with a Laplace approximation.
We compare LoID to the following baselines:
  • OOD-LR: Logistic regression model trained on OOD train set.
  • AutoElicit: LLM-generated expert-style priors via sampled completions.
@@
  • Cap: An oracle-style upper bound achieved by training a logistic regression model on the entire dataset, simulating access to the full population.
We evaluate each method using AUC score obtained from training on the ood train set and testing on the entire dataset.
5 Results
Table 2 summarizes the performance of each method across ten tabular datasets. Based on our experiments in Section A.2.2, Gemma-2-27B achieves the highest cumulative AUC among the evaluated LLMs. Therefore, Table 2 reports results obtained using this model.
5.1 AUC Comparison
LoID outperforms logistic regression trained on the OOD training set for all datasets except Liver Disease. Notably, on the Diabetes and Pima Diabetes datasets, LoID almost completely closes the performance gap induced by OOD training, indicating that it can successfully infer the underlying data distribution. On average, LoID closes 59% of the performance gap introduced by the OOD training setting.
LLMProcesses often underperforms even compared to uninformative priors. Since this approach does not explicitly produce priors, it is difficult to determine the exact cause of its poor performance. One possible explanation is its complete reliance on LLMs without leveraging traditional machine learning models, which are often more robust at pattern extraction in structured data.
LoID also achieves superior performance on 8 out of 10 datasets, achieving an average improvement of 99.7% in gap closed when it outperforms AutoElicit. To further understand these performance differences, we analyze the elicited prior distributions produced by both methods across selected winning and losing datasets. Section A.5.1 presents this comparison.
Finding 1: Non-Monotonic Relationship Between Prior Uncertainty and Performance
  • Optimal range ($\sigma\approx 0.4$$0.5$): Achieves best performance on 7 datasets: Liver, Income, Bank, Creditg, Diabetes, Pima, Occupancy
  • Too weak ($\sigma\gtrsim 0.8$): Underperforms AutoElicit: Blood, Heart
This shows that moderately weak priors ($\sigma\approx 0.4$$0.5$) provide sufficient regularization while allowing the Bayesian posterior to correct systematic biases in the elicited means through the data likelihood. Excessively weak priors ($\sigma\gtrsim 0.8$) fail to leverage LLM knowledge effectively, approaching uninformed priors.
LoID excels when Prior uncertainty is moderate, datasets have sufficient samples to correct sign biases, and AutoElicit’s tight priors impose incorrect strong beliefs.
Finding 3: Superior Mean Estimation Despite Directional Bias
While LoID exhibits systematic positive bias in sign coverage (0–6% negative priors vs. AutoElicit’s 20–50%), the absolute quality of the mean estimates appears superior. Evidence from performance gaps suggests:
@@ 5.1 AUC Comparison
    Magnitude accuracy over sign accuracy: On datasets where LoID wins, the improvement over AutoElicit indicates that the LLM-elicited means—despite sign bias—capture feature importance better than AutoElicit’s conversational approach.
@@
    Comparison to uninformed priors: LoID achieves  145% better gap closed vs. uninformed priors (see full comparison in A.6.1), demonstrating that the mean estimates contain genuine signal beyond just the regularization effect of uncertainty.
  3. 3.
    Interaction with uncertainty: The combination of (a) approximately correct magnitudes in means and (b) moderate uncertainty allows the Bayesian posterior to correct sign errors while preserving magnitude information. In contrast, AutoElicit’s tight uncertainties amplify errors in both magnitude and direction.
5.2 Computational Efficiency.
[…]
6 Discussion
The success of our LoID method challenges the conventional assumption that more accurate prior means necessarily lead to better performance. AutoElicit provides strong, confident priors ($\sigma\approx 0.05$$0.12$) with $\sim$40–50% correct negative signs. However, the incorrect confident priors harm inference by overpowering the data likelihood.
On the other hand, LoID provides weaker and uncertain priors ($\sigma\approx 0.4$$0.5$). This allows the weaker means to be overridden by data, providing soft regularization: the priors gently guide inference toward reasonable parameter spaces without rigidly enforcing incorrect beliefs. The result is that even when LoID’s directional beliefs are wrong, the posterior can recover because the prior is not insisting on its incorrectness. However, the benefits of this approach diminish when priors become too weak ($\sigma>0.8$).
@@ 7 Limitations and Future Work
Semantic granularity.
Our current approach probes binary directional relationships (positive vs. negative). Extending this framework to multi-token or scalar outputs could capture more nuanced associations, such as non-monotonic effects.
Hyperparameter selection:
Although we use fixed hyperparameters ($\alpha=0.2$, $\gamma=2.0$) across all test datasets, optimal values vary by task. Developing automated calibration methods or meta-learning approaches for selecting these parameters would improve out-of-the-box applicability.
LLM access requirements:
@@ 8 Conclusion
In this work, we introduce LoID, a deterministic method for extracting priors from LLMs. We show that priors elicited with LoID substantially mitigate the performance degradation caused by training on small, out-of-distribution datasets, recovering over 59% of the lost performance.
Our comparative analysis with AutoElicit demonstrates that the effectiveness of the LoID approach stems from honest uncertainty quantification, and correct mean magnitude estimation for logistic regression coefficients. By maintaining moderate prior uncertainty ($\sigma\approx 0.4$$0.5$), LoID provides soft regularization while allowing the Bayesian posterior to correct systematic biases through the data likelihood.
As a result, LoID outperforms AutoElicit on 80% of benchmark datasets, achieving an average improvement of 99.7% in gap closed when it wins. Against LLMProcesses, LoID achieves superior performance on all 10 datasets, with particularly dramatic gains on datasets where LLMProcesses’s end-to-end LLM approach fails to leverage structured data patterns.
@@ Appendix A Appendix
A.1 Reproducibility
Code Availability
Data Availability
All datasets used in this study are publicly available through open-access repositories. Further details and references are provided in Section A.4. All preprocessing steps required to obtain the final train/test splits used in our experiments, along with the preprocessed data, are available in our code release.
Results Variability
@@ A.2 Hyperparameter Optimization
We used three datasets—Blood Donation, Heart Disease, and Diabetes—as development datasets to tune hyperparameters based on AUC score.
@@ A.2.1 Number of Prompt Templates
We tested our method with 1–20 prompt templates analyzed how the number of templates affects LoID’s performance. The following are some examples of our prompt templates:
@@
As shown in Figure 2, performance gains plateau after approximately 10 sentences. Around this point, performance also becomes more stable, while incorporating additional sentences incurs increased computational cost. Accordingly, we use 10 sentences in all experiments.
Figure 2: Hyperparameter ablation results demonstrating that prior uncertainty (beta) is the critical factor for performance, with F-statistics of 346.7 (blood), 47.3 (heart), and 583.2 (diabetes), while variance sensitivity (alpha) and number of sentence formulations (num_sentences) show no significant effect.
@@ A.2.2 Model Selection
We compared the performances of LLMs–GPT-2, LLaMA 3.1 8B, Gemma-3-27B, and Gemma-2-27B in terms of AUC. The best averge performance was achieved with Gemma-2-27B, as shown in Table 3.
dataset blood diabetes heart average
model
gemma-3 0.65 0.84 0.89 0.79
gemma-2 0.67 0.84 0.90 0.80
llama-3-8b 0.56 0.80 0.85 0.73
gpt-2 0.53 0.82 0.84 0.71
Table 3: Model comparison across datasets using $\alpha=0.2$, $\gamma=2.0$, and 10 sentence formulations.
@@ A.2.3 Variance Estimation Formula
The prior standard deviation $\sigma_{j}$ for feature $j$ quantifies the uncertainty in the elicited prior mean. We explored two methods for computing $\sigma_{j}$ from LLM token probabilities, each capturing different aspects of elicitation uncertainty.
Method 1: Variance of Logit Differences
This method uses the variance across multiple sentence formulations as a measure of prior uncertainty:
$\sigma_{j}^{2}=\alpha+\gamma\cdot\text{std}(logit_{1}(p_{j}),\ldots,logit_{|T|}(p_{j}))$
@@
Rationale: Higher variance across sentence formulations indicates inconsistent LLM responses, suggesting greater epistemic uncertainty. The hyperparameters $\alpha$ and $\gamma$ control the base uncertainty level and sensitivity to variance, respectively.
@@
This method uses Shannon entropy to measure the uncertainty in the LLM’s token probability distribution:
$\sigma_{j}=\beta\cdot\frac{1}{N_{\text{sent}}}\sum_{i=1}^{N_{\text{sent}}}H(p_{\text{pos}}^{(i)},p_{\text{neg}}^{(i)})$
Rationale: Higher entropy (closer to $\log 2\approx 0.693$) indicates the LLM is uncertain between “positive” and “negative” tokens, reflecting ambiguity about the feature’s effect direction. Lower entropy indicates confident predictions. The scaling factor $\beta$ maps entropy to prior standard deviation.
Results
Table 4 compares the variance estimation methods described above. While the overall differences in performance are small, the variance of logit differences consistently performs slightly better across all three datasets. Accordingly, we adopt this method for LoID.
Dataset variance Entropy $\Delta$
Blood 0.67 0.66 +0.01
Heart 0.90 0.89 +0.01
Diabetes 0.84 0.83 +0.01
Average 0.80 0.79 +0.01
Table 4: Comparison of variance of logit differences (used by LoID) vs. entropy-based uncertainty for uncertainty estimation.
A.2.4 Tuning $\alpha$, $\gamma$, and $N_{\text{sent}}$
As detailed in Section 3.3, we use the following formula to calculate standard deviation:
$\sigma_{j}^{2}=\alpha+\gamma\cdot\text{std}(logit_{1}(p_{j}),\ldots,logit_{|T|}(p_{j}))$
We searched over $\alpha\in\{0.1,0.2,0.3,0.5\}$, $\gamma\in\{1,2,3,4\}$, and $N_{\text{sent}}\in\{5,10,15,20\}$. Table 5 reports the hyperparameter settings that achieve the best performance for each evaluation dataset. As shown in the table, no single hyperparameter configuration generalizes across all datasets; instead, each dataset requires task-specific tuning to achieve optimal performance. Among the evaluated $\alpha$$\gamma$ combinations, $\alpha=0.2$, $\gamma=2.0$, and $N_{\text{sent}}=10$ yield the best average performance. For simplicity, we adopt these values for all remaining experiments.
Dataset Best AUC $\alpha$ $\beta$ $N_{\text{sent}}$
Blood 0.6704 0.2 4.0 10
Heart 0.8987 0.1 2.0 20
Diabetes 0.8392 0.5 1.0 5
Overall (avg) 0.8027 0.2 2.0 10
Table 5: Best hyperparameter configurations identified for each dataset in the ablation study.
Figure 2 illustrates the relationship between AUC and the hyperparameters $\alpha$, $\gamma$, and the number of sentences. Among these, $\gamma$ has the most pronounced effect on performance, with its influence on AUC appearing approximately linear. For the Blood Donation dataset, AUC increases as $\gamma$ increases, whereas for the Heart Disease and Diabetes datasets, AUC decreases with larger values of $\gamma$. These trends indicate that prior certainty plays a pivotal role in model performance. In contrast, we observe no consistent relationship between $\alpha$, the number of sentences, and AUC across datasets, suggesting that these factors have a limited impact on performance.
Consistent with the previous findings, analysis of variance (ANOVA) reveals that $\gamma$ is the only statistically significant predictor of out-of-distribution performance across all datasets (Table 6). The baseline uncertainty parameter $\gamma$ exhibits strong effects ($p<0.0001$). In contrast, both $\alpha$ and $N_{\text{sent}}$ show no significant impact ($p>0.05$) on any dataset.
Dataset Parameter F-statistic p-value
Blood $\alpha$ 0.231 0.8743
$\gamma$ 346.743 <0.0001<0.0001
$N_{\text{sent}}$ 0.436 0.7280
Heart $\alpha$ 0.360 0.7820
$\gamma$ 47.269 <0.0001<0.0001
$N_{\text{sent}}$ 1.174 0.3285
Diabetes $\alpha$ 1.162 0.3331
$\gamma$ 583.151 <0.0001<0.0001
$N_{\text{sent}}$ 0.083 0.9691
Table 6: ANOVA results testing hyperparameter effects on out-of-distribution AUC. Only $\gamma$ (baseline prior uncertainty) shows significant effects across all datasets.

### v2

@@ 4.3 LLM-based prior extraction
We use Gemma-2-27B, Gemma-3-27B, llama-3-8B, and GPT-2 to extract logit-based priors as described in Section 3.2. We use these models in zero-shot mode without any fine-tuning or in-context examples. In our experiments, the queried lexical items are represented as single tokens across models, enabling direct logit-based comparisons.
Each dataset is processed independently. For each feature, we issue 10 prompts to extract token-level log-probabilities for the positive and negative completions. The number of prompts is selected via hyperparameter tuning (Section A.2.1). The resulting feature-wise prior means and variances are then used to define informative Gaussian priors in Bayesian logistic and linear regression models. Posterior inference is performed using PyMC with a Laplace approximation.
We compare LoID to the following baselines.
  • P_OOD: Logistic/Linear regression model trained on OOD train set.
  • AutoEl (AutoElicit): LLM-generated expert-style priors via sampled completions.
@@
  • P_IID: Performance obtained by training a logistic/linear regression model on the in-distribution (IID) data. This represents the oracle benchmark where the model has full access to the training distribution, serving as a reference point for evaluating OOD generalization.
  • N(0,1): An uninformed baseline using standard normal $\mathcal{N}(0,1)$ priors for all feature coefficients. This represents a weakly informative Bayesian approach without incorporating domain knowledge, serving as a reference point to quantify the information gain provided by LLM-derived priors. While a uniform prior would be maximally uninformative, our experiments in Section A.2.4 show that normal priors are more compatible with the MCMC pipeline and provide a stricter baseline.
For AutoElicit and LLMProcesses, we adopt the hyperparameters recommended by the original authors to the extent possible in order to ensure a fair comparison. A comprehensive list of the hyperparameters used to reproduce these baselines is provided in Section A.1.
5 Results and discussion
We evaluate LoID across 15 datasets spanning diverse domains to understand when LoID is able to close the performance gap caused by covariate shift. Our results in Table 1 reveal a systematic pattern: LoID excels when feature-target relationships are semantically meaningful and generalizable, but struggles when relationships are contextual, behavioral, or dataset-specific.
5.1 When does LoID help?
Medical and demographic tasks with stable relationships (83.8% gap closed on average). LoID achieves the strongest gains on datasets where feature-target relationships reflect generalizable biomedical or socioeconomic knowledge:
  • Adult Income (83.8% gap): Age, education, and occupation have well-known effects on income that generalize across demographic shifts
  • Concrete Strength (93.9% gap): Physical properties (cement, age, water) have deterministic relationships with compressive strength
  • Chronic Kidney Disease (46.2% gap): Clinical markers (hemoglobin, albumin) have established diagnostic value
In these domains, the LLM’s world knowledge captures stable causal mechanisms that remain valid under distribution shift.
Physical systems with known causal structure. For regression tasks where features have clear physical interpretations (Concrete Strength: 93.9% gap, Airfoil Noise: 68.9% gap, Communities and Crime: 62.5% gap), LoID leverages the LLM’s understanding of domain physics and social science to regularize predictions under shift.
Figure 2: LLM prior alignment across success and failure modes. Top: Well-aligned priors for Blood Donation and Chronic Kidney Disease show strong agreement between LLM beliefs and actual correlations. Bottom: Combined Cycle Power exhibits directionally correct but magnitude-misspecified priors, while Bank Marketing shows absent priors (near-zero) for context-dependent features.
5.2 When does LoID hurt?
Behavioral and contextual tasks with dataset-specific patterns. LoID fails on three datasets where relationships are inherently local or behavioral:
  • Bank Marketing ($130.6\%$ gap): Customer banking behavior (call duration, previous contacts) varies wildly across campaigns.
  • Bike Sharing ($1.8\%$ gap): Demand is driven by local weather, events, and infrastructure, relationships that are city-specific and temporal
  • Combined Cycle Power ($0.1\%$ gap): Plant efficiency depends on equipment-specific calibrations and operational conditions that LLMs cannot infer from generic turbine descriptions
In these cases, the LLM’s priors encode generic relationships (e.g., “higher temperature decreases power output”) that may be directionally correct but quantitatively misspecified for the specific plant, city, or campaign. This introduces harmful bias that degrades performance.
Understanding failure modes through prior analysis. Examining the extracted priors reveals two distinct failure modes:
@@ 5.2 When does LoID hurt?
    Overconfident but wrong priors destroy beneficial regularization. For Combined Cycle Power, the uninformed $\mathcal{N}(0,1)$ baseline closes 99.8% of the OOD gap (MSE: 23.58 vs. P_OOD: 962.94), demonstrating that simple Bayesian regularization is highly effective. However, LoID closes only 0.2% (MSE: 961.04), because the LLM’s equipment-agnostic priors (e.g., “higher ambient temperature decreases turbine efficiency”) override the beneficial regularization with domain knowledge that’s directionally plausible but quantitatively misspecified for this specific power plant and severe temperature shift (train: 1.8–13.5°C, test: 13.5–37.1°C). The LLM knows generic thermodynamics but cannot account for plant-specific calibrations or extreme extrapolation regimes.
@@
    Context-dependent features produce uninformative priors. For Bank Marketing and Bike Sharing, LLMs have no basis for judging how campaign-specific variables (e.g., “days since last contact”) or city-specific infrastructure affects outcomes. The extracted priors are either near-zero (indicating maximal uncertainty) or directionally inconsistent, providing no signal over uninformed regularization. Unlike Combined Cycle Power, where the problem is wrong priors, here the problem is absent priors—the LLM simply lacks relevant knowledge about these context-dependent domains.
Figure 2 illustrates four representative cases that reveal when and why LLM priors succeed or fail.
5.3 Comparison to baselines
LoID vs. LLMProcesses. LLMProcesses attempts to directly predict outcomes by prompting the LLM with feature values. While this works for more common tasks, it fails on real-world data: LLMProcesses achieves near-random performance (0.500 AUC) on 4/9 classification tasks (Stroke, Diabetes, Heart Failure, HNC). LoID outperforms LLMProcesses on 12/14 datasets (excluding Bank Marketing where both fail). This demonstrates that extracting priors is more effective than direct prediction—the LLM knows the relationships but cannot directly predict outcomes from sparse feature descriptions.
LoID vs. AutoElicit. AutoElicit elicits structured expert priors by prompting the LLM multiple times and aggregating responses to reduce variability. It performs well on tasks involving physical systems (Bike Sharing: 71405 MSE; Combined Cycle Power: 47.14 MSE), where aggregating numeric estimates helps calibrate prior ranges. However, for tasks requiring semantic reasoning about feature importance (Adult Income: 0.510 AUC; Concrete Strength: 1974.11 MSE), free-form elicitation may struggle to capture directional relationships as effectively as token-level probability extraction.
Moreover, LLMs tend to produce overconfident priors when explicitly prompted, which can limit the model’s ability to adapt to data. In contrast, LoID derives priors directly from token probabilities, leading to more calibrated uncertainty. As a result, LoID outperforms AutoElicit on 12/14 datasets.
The N(0,1) baseline reveals when priors help. Comparing LoID to the uninformed $\mathcal{N}(0,1)$ baseline isolates the value of LLM knowledge. On 11/15 tasks, LoID substantially outperforms $\mathcal{N}(0,1)$, confirming that LLM priors are informative.
@@ 6 Limitations and future work
  • LoID is most useful for dataset with meaningful semantic features, limiting its usage for de-identified datasets.
@@ 7 Conclusion
We introduced LoID, a method for extracting Bayesian priors from large language models through token-level probability distributions. By probing LLMs for beliefs about feature–target relationships and converting these into structured priors, LoID leverages embedded world knowledge to improve predictive performance under distribution shift.
Across 15 datasets, we observe a consistent pattern: LoID performs strongly when feature–target relationships are semantically stable and generalizable, closing 83.8% of the performance gap induced by out-of-distribution training in medical and demographic settings, but degrades when these relationships are context-dependent or dataset-specific, as in behavioral tasks such as Bank Marketing. Overall, LoID demonstrates that extracting distributional priors is more robust than direct prediction (compared to LLMProcesses) and more reliable than free-form elicitation (compared to AutoElicit), outperforming both baselines on 12 out of 15 datasets.
@@ Appendix A Appendix
A.1 Baseline hyperparameters
This section details the hyperparameters used for the baseline methods compared in our experiments. Table 2 provides a comprehensive overview of all configuration settings.
Category LLMProcesses AutoElicit LoID
Model Configuration
Model Gemma-2-27B-it Gemma-2-27B-it Gemma-2-27B-it
Temperature 1.0 1.0 N/A (no sampling)
Top-p 1.0 0.9 N/A
Max length 20 tokens 8192 tokens N/A
Prior/Prediction Extraction
Method Monte Carlo sampling Multi-expert prompting Token probability
Samples per point 50 500 (prior predictive) 10 templates
Aggregation Median MCMC mixture Mean and Variance of log-probs
Templates/Roles N/A 100 (10×10 role pairs) 10 sentence templates
Data Configuration
Training samples 50 (30 for HNC) 25 N/A (no in-context)
Prompt ordering Distance-based N/A N/A
Reproducibility
Random seed N/A 42 42
Repetitions 5 5 5
Table 2: Hyperparameter configurations for AutoElicit, LLMProcesses, and LoID.
@@ A.2 Hyperparameter optimization
We used four datasets—Stroke, Blood Donation, Hear Failure, and Adult Income—as development datasets to tune hyperparameters based on AUC and MSE.
@@ A.2.1 Number of prompt templates
We tested our method with 2–20 prompt templates and analyzed how the number of templates affects LoID’s performance. The following are some examples of our prompt templates:
@@
Figure 3 reveals the optimal template count for prior extraction. Performance peaks at 10 templates across most datasets, balancing signal aggregation against overfitting. Beyond 10 templates, performance degrades sharply (e.g., Stroke drops from 0.69 to 0.39 at 15 templates), suggesting that excessive prompting amplifies misspecified priors rather than refining them. While coefficient of variation decreases with more templates (indicating more stable uncertainty estimates), this stability comes at the cost of performance when templates exceed 10. Notably, heterogeneity reduction plateaus after 10 templates, yet performance continues to degrade, demonstrating that template averaging helps only when the underlying priors are well-specified. For datasets where LLM knowledge is weak (e.g., Blood), even 2 templates outperform 15+, as additional prompting merely reinforces uninformative/incorrect beliefs. This pattern supports our recommendation of 10 templates as the optimal trade-off between variance reduction and avoiding overconfident misspecification.
Figure 3: Template count sensitivity analysis across four datasets. Top two rows: Performance (Test AUC) versus number of sentence templates, comparing LoID to out-of-distribution baseline (P_OOD). Performance peaks at 10 templates for most datasets, with degradation at 15+ templates suggesting overfitting to misspecified priors. Bottom two rows: Prior heterogeneity measured by coefficient of variation (CV) of extracted prior standard deviations. Higher CV indicates greater variance in uncertainty across features. Heterogeneity decreases with more templates as averaging stabilizes estimates, but excessive templates (15+) harm performance without further reducing uncertainty.
@@ A.2.2 Model selection
We compared the performances of LLMs–GPT-2 (GPT), LLaMA-3.1-8B (L-3), Gemma-3-27B-Instruct (G-3), and Gemma-2-27B-Instruct (G-2) in terms of AUC. The best average performance was achieved with Gemma-2-27B, as shown in Table 3.
Dataset G-2 G-3 L-3 GPT-2
Stroke 0.694 0.690 0.624 0.712
Blood 0.663 0.522 0.609 0.579
Heart Failure 0.753 0.767 0.737 0.585
Adult 0.745 0.562 0.550 0.370
Average 0.714 0.635 0.630 0.562
(a) Model ablation.
Dataset M-1 M-2 M-3
Stroke 0.694 0.686 0.690
Blood 0.663 0.576 0.575
Heart Failure 0.753 0.708 0.743
Adult 0.745 0.674 0.659
Average 0.714 0.661 0.667
(b) Variance method ablation using Gemma-2-27B.
Table 3: Choosing the best model and variance estimate using four development datasets (Stroke, Blood Donation, Heart Failure, Adult Income).
@@ A.2.3 Variance estimation formula
The prior standard deviation $\sigma_{j}$ for feature $j$ quantifies the uncertainty in the elicited prior mean. We explored two methods for computing $\sigma_{j}$ from LLM token probabilities, each capturing different aspects of elicitation uncertainty. The best average performance was achieved with Method 1, as shown in Table 3.
Method 1: variance of normalized logit probabilities
In this method, $p_{j}$ reflects the model’s preference between logits. Applying a log transform maps probabilities from $[0,1]$ to the real line, aligning with the unbounded nature of regression coefficients. Variance across paraphrased inputs captures inconsistency in LLM responses, indicating greater epistemic uncertainty.
$\sigma_{j}^{2}=\text{std}(logit_{1}(p_{j}),\ldots,logit_{|T|}(p_{j}))$
@@
This method uses Shannon entropy to measure the uncertainty in the LLM’s token probability distribution. Higher entropy (closer to $\log 2\approx 0.693$) indicates the LLM is uncertain between the positive and negative tokens, reflecting ambiguity about the feature’s effect direction. Lower entropy indicates confident predictions.
$\sigma_{j}=\frac{1}{N_{\text{sent}}}\sum_{i=1}^{N_{\text{sent}}}H({P_{j}^{-}},P_{j}^{+})$
Method 3: standard deviation of log-probability differences
This method uses the difference of log-probabilities (equivalent to the log-ratio $\log(\frac{p_{\text{pos}}}{p_{\text{neg}}})$) without normalization, preserving information about the absolute probability mass assigned to both tokens. When templates produce inconsistent log-probability differences, this indicates epistemic uncertainty about both the direction and confidence of the feature’s effect. Unlike the normalized logit approach, this method is sensitive to cases where the LLM assigns low absolute probability to both tokens (indicating high uncertainty) versus high probability to one token (indicating confidence).
$\sigma_{j}^{2}=\text{std}(diff_{1}(p_{j}),\ldots,diff_{|T|}(p_{j}))$
Where
$diff(p_{j})=[\log(P_{j}^{+})-\log(P_{j}^{-})]$
A.2.4 Uninformative baseline: $\mathcal{N}$(0, 1) vs. $\mathcal{U}$(-1, 1)
We include an additional $\mathcal{N}(0,1)$ baseline to illustrate the improvement achieved over a more weakly informative prior that is commonly used with Bayesian inference.
Table 4 reveals that while Uniform priors ostensibly provide “maximum uninformativeness” by assigning equal probability across a range, they fail as regularizers for regression coefficients because:
  1. 1.
  2. 2.
    Incompatibility with MCMC: The flat density of Uniform priors provides no gradient information to guide sampling algorithms, resulting in inefficient exploration and convergence issues.
Dataset N(0,1) U(-1,1) $\Delta$
Stroke 0.694 0.448 -0.246
Blood 0.583 0.575 -0.008
Heart Failure 0.752 0.724 -0.028
Adult 0.745 0.761 +0.016
Mean 0.694 0.627 -0.066
Table 4: Comparison of Normal vs. Uniform uninformed priors for Bayesian logistic/linear regression.